Find the power necessary to move a 28.5 kg block up a frictionless 32.5 degrees incline at 4.65 m/s.
1
Expert's answer
2020-11-23T05:29:54-0500


In order to move a block up a frictionless incline one should apply force "\mathbf{F}" against the projection of the gravity force "m\mathbf{g}" on the x-axis. Here "m = 28.5kg" is the mass of the block and "g = 9.81m\/s^2" is the gravitational acceleration.

In order to maintain constant speed "v = 4.65m\/s" , this force should be:


"F = mg\sin32.5\degree"

The power is, then:


"P = Fv = mgv\sin32.5\degree\\
P = 28.5\cdot 9.81\cdot 4.65\cdot \sin32.5\degree \approx 698.5\space W"

Answer. 698.5 W.